OpenStudy (anonymous):

Rationalize the denominator of 6/(7-3 sqrt6)+(-2+2 sqrt6)

OpenStudy (anonymous):

@NewOne

OpenStudy (anonymous):

Is it \[\frac{ 6 }{ 7-3\sqrt{6} }+\left( -2+2\sqrt{6} \right)\]

OpenStudy (anonymous):

it's all under the 6

OpenStudy (anonymous):

sorry. I need to get my math done by tomorrow and I'm worrying haha

OpenStudy (anonymous):

Ok what you do first is simplify the denominator as much as you can.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

So what do you have as the denominator?

OpenStudy (anonymous):

6/5-1sqrt6?

OpenStudy (anonymous):

Yup, that's correct. Have you learned conjugates in class?

OpenStudy (anonymous):

yes! alright so times by (5+1sqrt6)?

OpenStudy (anonymous):

Multiply it into the denominator and numerator.

OpenStudy (anonymous):

Then you're done

OpenStudy (anonymous):

30+sqrt6/25+sqrt 6...?

OpenStudy (anonymous):

30-6sqrt6/31?

OpenStudy (anonymous):

Uh.. no Let's rewrite it so it's more clearer. \[\frac{ 6 }{ 5-1\sqrt{6}}*\frac{ 5+1\sqrt{6} }{ 5+1\sqrt{6} }\]

OpenStudy (anonymous):

Notice you're basically multiplying the fraction by 1

OpenStudy (anonymous):

so 30+6√6

OpenStudy (anonymous):

is the top

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

/19

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Whenever you are asked to rationalize, just remember to multiple the conjugate of the denominator to the top and bottom of the fraction